A pension fund manager is considering three mutual funds. The first is a stock fund, the second is a long-term government and corporate bond fund, and the third is a T-bill money market fund that yields a sure rate of 4.5%. The probability distributions of the risky funds are:   

Expected Return Standard Deviation
Stock fund (S) 15 % 35 %
Bond fund (B) 6 % 29 %

The correlation between the fund returns is .0517.


Suppose now that your portfolio must yield an expected return of 13% and be efficient, that is, on the best feasible CAL.


a. What is the standard deviation of your portfolio? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b-1. What is the proportion invested in the T-bill fund? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b-2. What is the proportion invested in each of the two risky funds? (Do not round intermediate calculations. Round your answers to 2 decimal places.)
Answer #1
To find the fraction of wealth to invest in Stock fund that will result in the risky portfolio with maximum Sharpe ratio the following formula to determine the weight of Stock fund in risky portfolio should be used

(E[Rp) – Rf Var(RE) - (E[RE] - RfCov(Re, Rp) (E[Rp) – Rf)Var(RE) + (E[RE] – Rp)Var(Rp) – (E[Rp] + E[RE] – 2R;)Cov(Rp, Rp)

Where
Stock fund E[R(d)]= 15.00%
Bond fund E[R(e)]= 6.00%
Stock fund Stdev[R(d)]= 35.00%
Bond fund Stdev[R(e)]= 29.00%
Var[R(d)]= 0.12250
Var[R(e)]= 0.08410
T bill Rf= 4.50%
Correl Corr(Re,Rd)= 0.0517
Covar Cov(Re,Rd)= 0.0052
Stock fund Therefore W(*d)= 0.8718
Bond fund W(*e)=(1-W(*d))= 0.1282
Expected return of risky portfolio= 13.85%
Risky portfolio std dev= 30.93%
Where
Var = std dev^2
Covariance = Correlation* Std dev (r)*Std dev (d)
Expected return of the risky portfolio = E[R(d)]*W(*d)+E[R(e)]*W(*e)
Risky portfolio standard deviation =( w2A*σ2(RA)+w2B*σ2(RB)+2*(wA)*(wB)*Cor(RA,RB)*σ(RA)*σ(RB))^0.5
Desired return = tbill return*proportion invested in tbill+risky portfolio return *proportion invested in risky portfolio
= tbill return*proportion invested in tbill+risky portfolio return *(1-proportion invested in tbill)
0.13=0.045*Proportion invested in Tbill+0.1385*(1-Proportion invested in Tbill)
Proportion invested in Tbill (answer b-1) = (0.1385-0.13)/(0.1385-0.045)
=0.0909 = 9.09%
proportion invested in risky portfolio = 1-proportion invested in tbill
=0.9091
Proportion invested in Bond fund (answer b-2) =proportion invested in risky portfolio *weight of Bond fund
=0.1165 = 11.65%
Proportion invested in Stock fund (answer b-2) =proportion invested in risky portfolio *weight of Stock fund
=0.7926 = 79.26%
std dev of portfolio (answer a) = std of risky portfolio*proportion invested in risky portfolio
0.9091*0.3093=28.12%